# Semi-Supervised Synthetic Data Generation with Fine-Grained Relevance Control for Short Video Search Relevance Modeling

### v1

@@ Introduction
Embedding models are fundamental components in many natural language processing applications, including web search, recommendation systems and so on(Huang et al. 2020; Zhao et al. 2023; Xi et al. 2025). With the rapid advancement of large language models (LLMs) and their remarkable performance across diverse tasks, recent research has explored leveraging LLMs to enhance embedding quality(Nie et al. 2024). A particularly promising direction involves synthesizing diverse and high-quality training data using LLMs, thereby improving the capability of embedding models to produce more accurate and semantically meaningful embeddings. Gemini Embedding (Lee et al. 2025) propose a multi-stage prompting strategy to synthesize both retrieval and classification datasets to improve data quality; Qwen3 embedding (Zhang et al. 2025) and KaLM-Embedding (Hu et al. 2025) leverage Persona Hub (Ge et al. 2024) to guide the data synthesis process, enabling the LLM to generate data from a variety of user personas to construct diverse synthetic train set.
Despite the effectiveness of these methods, existing LLM-based data synthesis methods are predominantly prompt-based and inherently limited by the generative capabilities of the LLM itself. As a result, there remains a noticeable gap between synthetic data and human-curated ground truth data—particularly in low-source domain-specific scenarios such as short video platforms (Zhang et al. 2024). Tang and Yang (2024) proposed FinMTEB, a counterpart to MTEB that consists of financial domain-specific text datasets, and observed that the state-of-the-art embedding models on MTEB got a significant performance drop on FinMTEB compared to their performance on MTEB. This highlights the necessity of synthesizing data that aligns with domain-specific distributions.
Moreover, current synthesis methods primarily focus on diversity in terms of attributes such as input length, difficulty, and user personas. Little attention has been paid to diversity in terms of semantic relevance, which plays a crucial role in downstream tasks like retrieval and ranking. Most synthesis methods adopt a binary relevance scheme (i.e., relevant vs. irrelevant) and rely on contrastive learning to train embedding models. Although hard negative mining(Zhang et al. 2022) has been widely employed to enhance the model’s sensitivity to subtle relevance differences, the binary formulation introduces a misalignment between the training objective and downstream task requirements. For example, retrieval tasks often demand fine-grained relevance ranking of documents rather than coarse binary relevance classification.
@@
We empirically validate SSRA by augmenting training data for the Qwen3-Embedding 4B model. Compared to the baseline trained on annotated data only, our method achieves a 1.73% improvement in nDCG@10 on the retrieval task and a 2.80% gain in average precision (AP) on pair classification. SSRA outperforms prompt-based synthetic data generation and vanilla supervised fine-tuning (SFT) synthetic data generation baselines, demonstrating its effectiveness in synthesizing domain-relevant data with target fine-grained relevance label. Moreover, our analysis highlights the importance of fine-grained semantic relevance diversity in improving downstream performance of embedding models. The embedding model trained with SSRA-augmented data has been deployed in the search enhanced recommendation pipeline of Douyin’s dual-column scenario, where it achieved significant gains in online A/B testing: CTR +1.45%, SRR +4.9%, and IUPR +0.1054%.
@@
Embedding Model Training and Evaluation.
@@
Embedding models serve as fundamental building blocks for a wide range of natural language processing (NLP) tasks, such as semantic similarity, information retrieval, and classification(Chandrasekaran and Mago 2021; Huang et al. 2020; Keraghel, Morbieu, and Nadif 2024). With the advent of pre-trained language models(PLMs), the learning paradigm of embedding models gradually unify to contrastive learning(Wu et al. 2020; Gao, Yao, and Chen 2021; Izacard et al. 2021). Contrastive learning focuses on constructing positive and negative sample pairs, and optimizes the embedding space by pulling positive pairs closer and pushing negative pairs farther apart. A typical loss function used in this framework is the InfoNCE loss(Oord, Li, and Vinyals 2018). Recent state-of-the-art LLM-based embedding models(Lee et al. 2025; Zhang et al. 2025), whose backbone are LLMs(Team et al. 2023, 2024; Yang et al. 2025), continue to follow the contrastive learning paradigm, focusing on enhancing model performance through higher-quality positive and negative pairs, as well as improved loss functions based on InfoNCE loss.
The advancement of embedding models also relies heavily on comprehensive and standardized evaluation benchmarks. To address this need, the Massive Text Embedding Benchmark (MTEB) was introduced to provide a unified framework for embedding evaluation (Muennighoff et al. 2022; Enevoldsen et al. 2025). The MTEB benchmark contains 56 datasets in 7 different types of downstream tasks, providing an efficient and comprehensive evaluation framework for embedding models.
Synthetic Data Generation for Embedding Models.
@@
Synthetic query generation (Liang et al. 2020; Jeronymo et al. 2023) based on given documents has been widely used to create diverse training data for embedding models. Lee et al. (2024) showed that documents used to synthesize queries may be suboptimal as positive examples. To address this, their method retrieves candidate documents using the generated query and selects the most relevant one via LLM-based scoring. In addition, many studies directly synthesize positive and hard negative documents as training data for embedding-based retriever tuning. Kim and Baek (2025) leveraged LLMs to synthesize both positive and hard negative documents, aiming to distill the natural language understanding capabilities of LLMs into a lightweight retriever. Shao et al. (2025) proposed a data synthesis pipeline that generates a challenging yet relevant query and a corresponding hard negative given a document, facilitating the construction of high-quality training triplets.
@@
While these work improved embedding models by synthesizing high-quality and diverse training data, the notion of relevance is still limited to binary labels. This binary supervision introduces a fundamental discrepancy between the training objective of embedding models and the requirements of downstream tasks such as retrieval and reranking. To overcome this limitation, Esfandiarpoor et al. (2025) proposed a synthesis approach to generate 4-level relevance documents for a given query using LLMs. However, their data synthesis approach remains prompt-based and is inherently constrained by the capabilities of existing LLMs, particularly when generating data for specific domain at smaller scales—as acknowledged in their own limitations. Moreover, their definition of relevance is generic and simplistic, overlooking the domain-specific variations in how relevance should be modeled and interpreted.
@@
Step 1: Query-Item Collection. We first collect query-item pairs, where each item refers to a short video on the short video platform. To obtain high-quality query-item candidates for annotation, we leverage search click logs of a major Chinese short video platform. Specifically, we employ two complementary strategies:
@@
Method Qwen3-embedding 0.6B Qwen3-embedding 4B
@@
Table 4: Performance of models trained with SSRA augmented data with binary(0,1) and multi-level(0,1,2,3) relevance labels. Note that the AP results are calculated in a balanced-version pair classification test set.
@@
To train both the query model and the score model, we adopt Doubao-1.5-Pro-32K as the backbone and both models are trained using a batch size of 16 and an initial learning rate of $1\times 10^{-5}$. To utilize 4-level relevance labels, we utilize infoNCE loss with label as additional weight of loss.
Following the two-stage SSRA pipeline, we use the final query model to synthesize relevance-controlled queries for 1 million distinct short video documents, covering relevance labels of 1, 2, and 3444We use in-batch negative method to get negative pairs, whose relevance label is 0.. The synthesized queries are then filtered using the score model to ensure label consistency. These high-quality synthetic queries are combined with the proposed annotated train set to form the final train set for the embedding model.
We employ Qwen3-Embedding [^5] https://huggingface.co/Qwen as the backbone for embedding model training, and utilize both in-batch negative sampling and hard negative mining techniques. We train embedding models at two parameter scales: 0.6B and 4B. For both settings, we adopt the Qwen3-Embedding architecture with lora fine-tuning, following the implementation from the Qwen3-Embedding blog [^6] https://qwenlm.github.io/blog/qwen3-embedding/. Specifically, we apply LoRA (Hu et al. 2022) with a rank of 32 and the models are trained with a batch size of 512 and a learning rate of $1\times 10^{-5}$.
@@
For the retrieval test set, we follow the MTEB(Muennighoff et al. 2022) evaluation protocol and report nDCG@10 as the primary metric. For the pairwise classification test set, we also follow the MTEB standard and report Average Precision (AP) at different relevance thresholds (label $\geq$ 1, 2, and 3) to assess semantic matching quality at varying levels of granularity.
@@
We consider several baselines to evaluate the effectiveness of our SSRA pipeline. First, we include a Base Model trained directly on the annotated training data without any synthetic augmentation. Second, we evaluate a Vanilla SFT baseline, where synthetic data is generated by a query model trained on the annotated dataset and used to augment training dataset. Third, we include a prompt-based baseline, SyCL Modified, adapted from SyCL(Esfandiarpoor et al. 2025) which originally generates documents of varying relevance labels given a query. We retain its core prompt design but modify the objective to generate relevance-controlled queries given a document, making it a representative of prompt-based query synthesis methods.
@@
Method Qwen3-embedding 0.6B Qwen3-embedding 4B
@@
Table 5: Ablation results of the two-stage SSRA pipeline.
@@
We randomly sample 40 documents and use query generation models trained with only Stage 1 and with both Stage 1 and Stage 2 to generate queries intended to match relevance levels 1, 2, and 3, respectively. The generated query-document pairs are then annotated by human annotators using the same 4-level relevance schema. We record the number of generated queries whose annotated relevance label matches the intended label.
The consistency results are summarized in Table 6. As shown, the model trained with Stage 2 significantly improves the alignment between intended and actual relevance levels with a 24.29% improvement in consistency, demonstrating enhanced controllability and generation precision.
@@
Based on the embedding model trained with synthetic data from SSRA pipeline, we conducted rigorous online A/B tests in the dual-column scenario of a major Chinese short video platform from May 9 to May 19, 2025, with hundreds of millions of user participations per day. The embedding model is used to filter out videos irrelevant to the query, ensuring that the displayed videos are closely related to users’ historical search queries. The results are presented in Table 7, focusing on several key metrics: CTR (Click-Through Rate), SRR (Strong Relevance Ratio ), and IUPR (Image User Penetration Rate). SRR is defined as the percentage of content strongly relevant to the user’s search query among the videos ultimately recommended. IUPR refers to the proportion of users who clicked on image content among the total daily active users, which is a core metric in the dual-column scenario.
@@ Conclusion
For future work, we will explore task-specific loss functions tailored to fine-grained relevance labels in training data. Since different tasks (e.g., retrieval vs. pair classification) have distinct objectives, customized loss designs can better leverage supervision signals to maximize their respective performance.
@@ Appendix A Appendix
However, our annotated dataset only contains scalar relevance labels without intermediate reasoning steps. To address this, we leverage the DeepSeek-R1(Guo et al. 2025) model to generate full reasoning chains. Specifically, given a query-document pair and its associated relevance label, we apply the prompt design illustrated in Figure 4 to generate the missing analytical reasoning. This enables the construction of a training dataset tailored for two-stage relevance modeling.

### v2

@@ Introduction
Embedding models are fundamental components in many natural language processing applications, including web search, recommendation systems and so on(huang2020embedding; zhao2023embedding; xi2025aug2search). Recent studies have explored using LLMs to improve embedding quality (nie2024text), particularly by synthesizing diverse, high-quality training data to enhance the semantic accuracy of embeddings. Gemini Embedding (lee2025gemini) propose a multi-stage prompting strategy to synthesize both retrieval and classification datasets to improve data quality; Qwen3 embedding (zhang2025qwen3) and KaLM-Embedding (hu2025kalm) leverage Persona Hub (ge2024scaling) to guide the data synthesis process, enabling the LLM to generate data from a variety of user personas to construct diverse synthetic train set.
Despite the effectiveness of these methods, existing LLM-based data synthesis methods are predominantly prompt-based and inherently limited by the generative capabilities of the LLM itself. As a result, there remains a noticeable gap between synthetic data and human-curated ground truth data—particularly in low-source domain-specific scenarios such as short video platforms (zhang2024value). tang2024we proposed FinMTEB, a counterpart to MTEB (muennighoff2022mteb; enevoldsen2502mmteb) that consists of financial domain-specific text datasets, and observed that the state-of-the-art embedding models on MTEB got a significant performance drop on FinMTEB compared to their performance on MTEB. This highlights the necessity of synthesizing data that aligns with domain-specific distributions.
Moreover, current synthesis methods primarily focus on diversity in terms of attributes such as input length, difficulty, and user personas. Little attention has been paid to diversity in terms of semantic relevance, which plays a crucial role in downstream tasks like retrieval and ranking. Most synthesis methods adopt a binary relevance scheme (i.e., relevant vs. irrelevant) and rely on contrastive learning to train embedding models. Although hard negative mining(zhang2022unsupervised) has been widely employed to enhance the model’s sensitivity to subtle relevance differences, the binary formulation introduces a misalignment between the training objective and downstream task requirements. For example, retrieval tasks often demand fine-grained relevance ranking of documents rather than coarse binary relevance classification.
@@
We empirically validate SSRA by augmenting training data for the Qwen3-Embedding 4B model. Compared to the baseline trained on annotated data only, our method achieves a 1.73% improvement in nDCG@10 on the retrieval task and a 2.80% gain in average precision (AP) on pair classification. SSRA outperforms prompt-based synthetic data generation and vanilla supervised fine-tuning (SFT) synthetic data generation baselines, demonstrating its effectiveness in synthesizing domain-relevant data with target fine-grained relevance label. Moreover, our analysis highlights the importance of fine-grained semantic relevance diversity in improving downstream performance of embedding models. The embedding model trained with SSRA-augmented data has been deployed in the search enhanced recommendation pipeline of a Douyin’s dual-column scenario, where it achieved significant gains in online A/B testing: CTR +1.45%, SRR +4.9%, and IUPR +0.1054%.
@@
Embedding Model Training
@@
Embedding models serve as fundamental building blocks for a wide range of natural language processing (NLP) tasks, such as semantic similarity, information retrieval, and classification(chandrasekaran2021evolution; huang2020embedding; keraghel2024beyond). With the advent of pre-trained language models(PLMs), the learning paradigm of embedding models gradually unify to contrastive learning(wu2020clear; gao2021simcse; izacard2021unsupervised). Contrastive learning focuses on constructing positive and negative sample pairs, and optimizes the embedding space by pulling positive pairs closer and pushing negative pairs farther apart. A typical loss function used in this framework is the InfoNCE loss(oord2018representation). Recent state-of-the-art LLM-based embedding models(lee2025gemini; zhang2025qwen3), whose backbone are LLMs(team2023gemini; team2024gemini; yang2025qwen3), continue to follow the contrastive learning paradigm, focusing on enhancing model performance through higher-quality positive and negative pairs, as well as improved loss functions based on InfoNCE loss.
Synthetic Data Generation for Embedding Models
@@
Synthetic query generation (liang2020embedding; jeronymo2023inpars) based on given documents has been widely used to create diverse training data for embedding models. lee2024gecko showed that documents used to synthesize queries may be suboptimal as positive examples. To address this, their method retrieves candidate documents using the generated query and selects the most relevant one via LLM-based scoring. Moreover, many studies directly synthesize positive and hard negative documents as training data for embedding-based retriever tuning. kim2025syntriever leveraged LLMs to synthesize both positive and hard negative documents, aiming to distill the natural language understanding capabilities of LLMs into a lightweight retriever. shao2025reasonir proposed a data synthesis pipeline that generates a challenging yet relevant query and a corresponding hard negative document given a document, facilitating the construction of high-quality training triplets.
@@
While these work improved embedding models by synthesizing high-quality and diverse training data, the notion of relevance is still limited to binary labels. This binary supervision introduces a fundamental discrepancy between the training objective of embedding models and the requirements of downstream tasks such as retrieval and reranking. To overcome this limitation, esfandiarpoor2025beyond proposed SyCL, a prompt-based approach to generate 4-level relevance documents for a given query using LLMs.
@@
Step 1: Query-Item Collection. We first collect query-item pairs, where each item refers to a short video on the short video platform. To obtain high-quality query-item candidates for annotation, we leverage search click logs of Douyin. Specifically, we employ two complementary strategies:
@@
Method 0.6B Model 4B Model
@@
Table 4: Performance of Qwen3-embedding trained with SSRA augmented data with binary(0,1) and multi-level(0,1,2,3) relevance labels. Note that the AP results are calculated in a balanced-version pair classification test set.
@@
[…]
Following the two-stage SSRA pipeline, we use the final query model to synthesize relevance-controlled queries for 1 million distinct short video documents, covering relevance labels of 1, 2, and 3555We use in-batch negative method to get negative pairs, whose relevance label is 0.. The synthesized queries are then filtered using the score model to ensure label consistency. These high-quality synthetic queries are combined with the proposed annotated train set to form the final train set for the embedding model.
We employ Qwen3-Embedding [^6] https://huggingface.co/Qwen as the backbone for embedding model training, and utilize both in-batch negative sampling and hard negative mining techniques. For We train embedding models at two parameter scales: 0.6B and 4B. For both settings, we adopt the Qwen3-Embedding architecture with lora fine-tuning, following the implementation from the Qwen3-Embedding blog [^7] https://qwenlm.github.io/blog/qwen3-embedding/. Specifically, we apply LoRA (hu2022lora) with a rank of 32 and the models are trained with a batch size of 512 and a learning rate of $1\times 10^{-5}$.
@@
For the retrieval test set, we follow the MTEB(muennighoff2022mteb) evaluation protocol and report nDCG@10 as the primary metric. For the pairwise classification test set, we also follow the MTEB standard and report Average Precision (AP) at different relevance thresholds (label $\geq$ 1, 2, and 3) to assess semantic matching quality at varying levels of granularity.
@@
We consider several baselines to evaluate the effectiveness of our SSRA pipeline. First, we include a Base Model trained directly on the annotated training data without any synthetic augmentation. Second, we evaluate a Vanilla SFT baseline, where synthetic data is generated by a query model trained on the annotated dataset and used to augment training dataset. Third, we include a prompt-based baseline, SyCL Modified, adapted from SyCL(esfandiarpoor2025beyond) which originally generates documents of varying relevance labels given a query. We retain its core prompt design but modify the objective to generate relevance-controlled queries given a document, making it a representative of prompt-based query synthesis methods.
@@
We find that, for the retrieval task, training data with intermediate relevance levels leads to a notable improvement in nDCG@10, highlighting the importance of fine-grained and diverse relevance labels in enhancing retrieval performance.
@@
Method 0.6B Model 4B Model
@@
Table 5: Ablation results of the two-stage SSRA pipeline using Qwen3-embedding.
@@
We randomly sample 200 documents and use query generation models trained with only Stage 1 and with both Stage 1 and Stage 2 to generate queries intended to match relevance levels 1, 2, and 3, respectively. The generated query-document pairs are then annotated by human annotators using the same 4-level relevance schema. We record the number of generated queries whose annotated relevance label matches the intended label.
As shown in Table 6, the model trained with Stage 2 significantly improves the alignment between intended and actual relevance levels with a 25.43% relative improvement in consistency, demonstrating enhanced controllability and generation precision.
@@
Based on the embedding model trained with synthetic data from SSRA pipeline, we conducted rigorous online A/B tests in the dual-column scenario of Douyin from May 9 to May 19, 2025, with hundreds of millions of users participating per day. Users were randomly assigned to control and experimental groups via hash bucketing (190 million users each). The embedding model is used to filter out videos irrelevant to the query, ensuring that the displayed videos are closely related to users’ historical search queries.
The results are presented in Table 7, focusing on several key metrics: CTR (Click-Through Rate), SRR (Strong Relevance Ratio), and IUPR (Image User Penetration Rate). SRR is defined as the percentage of content strongly relevant to the user’s search query among the videos ultimately recommended. IUPR refers to the proportion of users who clicked on image content among the total daily active users, which is a core metric in the dual-column scenario.
@@ Conclusion
Acknowledgments
We sincerely appreciate our colleagues at ByteDance for their support. They contributed to this work but were not listed as authors. All contributors are listed in alphabetical order by last name: Zirui Guo, Jianheng Ma, Yaling Mou, Lu Ruan, Hailiang Wang.
@@ Appendix A Appendix
However, our annotated dataset only contains scalar relevance labels without intermediate reasoning steps. To address this, we leverage the DeepSeek-R1(guo2025deepseek) model to generate full reasoning chains. Specifically, given a query-document pair and its associated relevance label, we apply the prompt design illustrated in Figure 4 to generate the missing analytical reasoning. This enables the construction of a training dataset tailored for two-stage relevance modeling.